Paula's current salary as a Systems Engineer is $120,000 per year. She's planning to place 15% of her salary each year in a mutual fund. Paula expects a $4000 salary increase each year for the next 25 years of employment. If the mutual fund will average 10% per year, what will be the amount of interest accumulated at the time of her retirement?
Answer #1

First investment = 15% * 120000 = 18000

Increase in investment every year = 4000 * 15% = 600

t= 25 yrs

i =10%

Future value of this investment = A *(F/A,10%,25) + G (F/G, 10%, 25)

= 18000 * 98.3470594 + 600* 733.470594

= 2,210,329.43